OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10

OpenStudy (anonymous):

kay all you have to do is just substitute one equation into the other. So just isolate any variable in one equation then substitute that into the second equation to solve for the other variable

OpenStudy (anonymous):

2y + 10?

OpenStudy (anonymous):

k so for this question isolate x or y in any equation. I choose to rearrange the second equation. Rearrange (2) -x - 2y = -10 -x = 2y - 10 * multiply both sides by -1 x = 10 - 2y substitute x = 10 - 2y into equation (1) and solve for y

OpenStudy (anonymous):

no i mean is 2y + 10 the answer to the question

OpenStudy (anonymous):

i dont get what ur saying...... answer to what question

OpenStudy (anonymous):

oh no the answer is wrong

OpenStudy (anonymous):

10 - 2y is what you would substitute

OpenStudy (anonymous):

so its -2y - 10?

OpenStudy (anonymous):

no its -2y + 10